There are 'm' points on a straight line AB and 'n' points on the line AC none of them being the point A. Triangles are formed with these points as vertices, when


 (i)A is excluded (ii) A is included. The ratio of no of triangles in the 2 cases is ? 

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear jee king

case 1   A is excluded

total ways =m+nC3

but triangle can't be formebd joining the 3 points which are in straight line so

total number of possible triangle=m+nC3  -mC3  -nC3

 

Case 2 A is included

otal ways =m+n+1C3

but triangle can't be formebd joining the 3 points which are in straight line so

total number of possible triangle=m+n+1C3  -m+1C3  -n+1C3


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin